Fermi National Laboratory
 

Operation Summary for May 28, 2010 - May 31, 2010

 

The Integrated Luminosity for the period from 5/24/10 to 5/31/10 was 36.88 pb-1.
NuMI reported receiving 8.37E18 protons on target during this same period.

Friday May 28

The day shift began with Operations monitoring the Recycler stash (0.9E10) and the Antiproton Source (Pbar) stack (26.4E10), with the Linac and the Main Injector (MI) operating normally, but with the Booster in access, which means no downstream beam, with the NuMI and MiniBooNE experiments waiting for beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) waiting for the next store.

During the day shift, the Booster access ended and the turn-on began at 11:45 AM. The Tevatron (TeV) remained in standby until noon at which point cryo system technicians addressed TeV sector C3 wet engine bearing issues. The 400 MeV area chopper developed a high voltage cable problem that delayed beam to the Pbar, NuMI, and MiniBooNE until 2:15 PM. The C3 cryo system resumed normal operations at 3 PM. TeV personnel completed a dry squeeze and then began a 150 GeV study. Operations eventually delivered beam to the Pbar Source, NuMI, MiniBooNE, and FTBF.

During the evening shift, TeV personnel completed their studies and then left the machine in a low beta squeeze. Operations put the TeV into shot setup at 8:05 PM and established store 7854 at 9:10 PM with an average initial luminosity (AIL) of 167.8E30. The antiproton transfer line (AP0) pressure and airflow came into alarm at 10:40 PM. Operations talked with Pbar and Safety personnel. Operations halted stacking and prepared AP0 for an access by Safety personnel.

Saturday May 29

The midnight shift began with Operations monitoring store 7854 (92.5E30), the stash (73.8E10), and the stack (14.5E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations reported that Pbar was off due to AP0 air handling system work. Operations established beam to the FTBF T-1006 experiment at 4:15 AM.

During the day shift, Pbar resumed stacking. Operations monitored systems.

During the evening shift, Operations disabled the T-1006 experiment at 6 PM. Operations terminated store 7854 at 6:40 PM and put the TeV into shot setup. Operations established store 7855 at 7:55 PM with an AIL of 296.07E30. Operations reported a chilled water system leak at AP50; FESS personnel repaired the leak, but the work interrupted stacking for ~1.5 hours. Operations reported some minor downtime when Booster beam valves closed for no apparent reason around 8:30 PM.

Sunday May 30

The midnight shift began with Operations monitoring the store 7855 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations established beam to T-1006 at 4:30 AM. At 5 AM, the Booster East Anode power supply tripped off on an AC input overload and on a tunnel LCW pressure indication; everything reset.

During the day shift, Operations terminated store 7855 at 12:38 PM and put the TeV into shot setup, but the store quenched at step 24 of the squeeze at TeV sectors DA, DB, DD, and D1. Operators investigated, but could not find any obvious cause for the quench. Operators prepared the CDF and D0 collision halls for accesses.

During the evening shift, CDF and D0 came out of access at 4:30 PM. The TeV quenched again at DA and DB during the first HEP ramp. A TeV person came in and another person helped from home with the investigation, which lasted throughout the shift. Operations reported problems with Booster ion pumps (IPRF14 and IP142) tripping off frequently; Operations informed Booster personnel. Operations disabled the T-1006 beam at 6 PM.

Monday May 31

The midnight shift began with Operations monitoring the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations reported that TeV personnel could not find a reason for either of the quenches. TeV personnel parsed the squeeze and made necessary adjustments. Operations put the TeV into shot setup at 12:18 AM and established store 7860 at 1:29 AM with an AIL of 277E30. IPRF14 tripped several times during the shift and caused some downtime. Operations established beam to T-1006 at 4:30 AM.

The Plans for Monday

The plans for today are to stack, stash, and store.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 06/01/2010    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory